OpenStudy (anonymous):

A man invests Rs 3,20,000 for 2 years at 12.5% p.a. compounded annually. If the income tax at the rate of 20% is deducted at the end of each year, on interest accrued, find the amount he received at the end of 2 years.

OpenStudy (anonymous):

320000*12.5% = 40000 40000*20% = 8000 40000-8000 = 32000 320000+32000= 352000 repeat for second year

OpenStudy (anonymous):

320000*12.5% *2= 8000000 8000000*20% = 1600000 8000000-1600000 = 6400000 6400000+6400000= 12800000

OpenStudy (anonymous):

Don't think that is right

OpenStudy (anonymous):

So 320000 is the starting value after one year 12.5% of the 320000 is added as interest, but before it is added it is reduced by the income tax by 20%. For the second year you start with 352000

OpenStudy (anonymous):

i didn't understand can u plse do it

OpenStudy (anonymous):

352000*12.5% = 44000 44000*20% = 8800 44000-8800 = 35200 352000+35200 = 387200

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

yes soo this the final answer

OpenStudy (anonymous):

It should be if I understood the question.